Vanillin, the primary flavoring component in vanilla, is a molecule that contains the elements carbon, hydrogen, and oxygen. Complete combustion of a 0.420 g sample of vanillin produced 0.972 g carbon dioxide and 0.199 g of water.
Find the empirical formula for vanillin
1
Expert's answer
2019-10-07T06:43:00-0400

nCO2 = 0.972/44 = 0.022

nC = 0.022

nH2O = 0.199/18= 0.011

nH = 0.011*2= 0.022

nO =( 0.420-(0.022*12+0.022*1)) /16=0.008

C0.022H0.022O0.008=C8H8O3